# Distributed Model Checking in Graphs Classes

### v1

@@ 1.1.2 Counting and Optimization
Let $\Lambda$ be a finite set. For every quantifier-free FO formula $\varphi(x_{1},\dots,x_{k})$ on $\Lambda$-labeled graphs with $k\geq 1$ free variables $x_{1},\dots,x_{k}$, and for every class of labeled graphs $\mathcal{G}$ of bounded expansion, there exists a distributed algorithm that, for every $n$-node network $G\in\mathcal{G}$ of diameter $D$ and every labeling function $L:V\to 2^{\Lambda}$, counts the number of $k$-tuples of vertices $(v_{1},\dots,v_{k})$ for which $G\models\varphi(v_{1},\dots,v_{k})$, in ${O(D+\log n)}$ rounds under the CONGEST model.
Optimization problems can also be expressed as fining optimal solutions for FO formulas with free variables, over weighted graphs, e.g., finding a maximum weight triangle. As usual in the literature on CONGEST [69], the weights assigned to the nodes or edges are assumed to be polynomial in the network size (so that any weight can be transmitted between neighbors in a single round of communication). We prove the following theorem.
Theorem 3.
@@
Let $\Lambda$ be a finite set. For every FO formula $\varphi(x_{1},\dots,x_{k})$ on $\Lambda$-labeled graphs with $k\geq 1$ free variables $x_{1},\dots,x_{k}$, and for every class of labeled graphs $\mathcal{G}$ of bounded expansion, there exists a distributed algorithm that, for every $n$-node network $G=(V,E)\in\mathcal{G}$ of diameter $D$, every labeling function $L:V\to 2^{\Lambda}$, and for every weight function $\omega:V\to\mathbb{N}$, computes a $k$-tuple of vertices $(v_{1},\dots,v_{k})$ such that $G\models\varphi(v_{1},\dots,v_{k})$ and $\sum_{i=1}^{k}\omega(v_{i})$ is maximum, in $\mathcal{O}(D+\log n)$ rounds under the CONGEST model. (The same holds if replacing maximum by minimum.)
@@
Note that, in contrast to our Theorem 2, the best known algorithm for counting triangles in general graphs takes $\tilde{\mathcal{O}}(\sqrt{n})$ rounds [9]. Theorem 3 allows to find, for example, the maximum weight of a triangle or of a $k$-independent set.
1.1.3 Distributed Certification
Last but not least, our results extend to distributed certification.
Theorem 4.
@@ 1.1.3 Distributed Certification
Let $\Lambda$ be a finite set. For every FO formula $\varphi$ on labeled graphs, and for every class of $\Lambda$-labeled graphs $\mathcal{G}$ of bounded expansion, there exists a distributed certification scheme that, for every $n$-node network $G=(V,E)\in\mathcal{G}$ and every labeling $L:V\rightarrow 2^{\Lambda}$, certifies $(G,L)\models\varphi$ using certificates on $\mathcal{O}(\log n)$ bits.
Note that, with $O(\log n)$-bit certificates, the verifier can be implemented in a single round of communication in the CONGEST model, as desired. Note also that our certification scheme is a so-called proof-labeling scheme [57], which is a weaker form of distributed certification than, e.g., locally checkable proofs [46].
@@
As a direct consequence of Theorem 4, certifying $C_{4}$-freeness can be done with $\mathcal{O}(\log n)$-bit certificates in graphs of bounded expansion. In contrast, certifying $C_{4}$-freeness requires certificates on $\tilde{\Omega}(\sqrt{n})$ bits in general graphs, using the reduction from [15]. More generally, [22] show that every MSO formula can be certified with $O(\log n)$-bit certificates in graphs of bounded treedepth. Theorem 4 extends this result. Indeed, while Theorem 4 is for FO, and the result of [22] is for MSO, both logics have identical expressive power on graphs of bounded treedepth.
Last but not least, as for our model-checking algorithm, Theorem 4 is the best that can be achieved. Indeed, it does not extend to MSO as certifying non-3-colorability requires certificates of $\tilde{\Omega}(\sqrt{n})$ bits in planar graphs (see Theorem 6), and certificates of $\tilde{\Theta}(n)$ bits in bounded degree graphs (see Theorem 5).
@@ 1.2 Our Techniques
1.2.1 Distributed Model Checking
The proof of our model checking algorithm in Theorem 1 uses tools completely different from the low-congestion short-cut technique developed for planar graphs and graphs of bounded genus by [39, 40]. It is also different from the ones in [28, 22, 32, 31] used for graphs of bounded treedepth, treewidth, and cliquewidth, respectively. Our motivation is to bring in the techniques and ideas developed for efficient centralized meta-algorithms for sparse classes of graphs to the distributed domain. In a nutshell, we use modern tools from centralized model checking, and we show how to implement them efficiently in the distributed setting, namely the CONGEST model. Our approach is based on the quantifier elimination methods from [16], [48], and, mainly, [71].
Before we summarize our approach, let us note that model checking FO formulas in graphs of bounded expansion is trivial in the so-called Congested Clique model (referred to as CLIQUE hereafter), which is essentially the CONGEST model in a complete graph $K_{n}$. In this model, every node $i\in\{1,\dots,n\}$ of $K_{n}$ plays the role of a node $v$ of the input graph $G=(V,E)$ (i.e., node $i$ is given as input all the edges $\{i,j\}$ incident to $v$ in $G$). Now, in the CLIQUE model, every graph of bounded degeneracy can be reconstructed at each node in a single round of communication (see [2]). This holds even in the broadcast variant BCLIQUE of CLIQUE in which every node is bounded to send a same $\mathcal{O}(\log n)$-bit message to all the other nodes at each round. The techniques used for obtaining this result in CLIQUE do not extend to CONGEST as, in the latter, nodes can only communicate through the edges of the input graph $G$, and thus every node may only learn $\mathcal{O}(\log n)$ bits of information from each of its neighbors at each round. Instead, in BCLIQUE, each node can learn a total of $\Theta(n\log n)$ bits of information at each round even if $G$ is a graph of bounded degree.
We implement in the CONGEST model a quantifier elimination technique designed for the centralized setting, so that the elimination is performed in $\mathcal{O}(D)$ rounds in diameter-$D$ graphs of bounded expansion. For this, we essentially revisit the existing approaches, and we show that they can be implemented in the distributed setting. As said before, our argument is strongly based on the exposition of [71], which in turn is based on the work of [48]. However, our work differs from the previous ones in two important aspects.
First, the distributed implementation of the techniques in [71] requires technical attention in several places. There are indeed operations that are very simple to perform in the centralized setting, and they are actually often even implicit when describing the sequential algorithm. We had to carefully revisit the whole construction of Pilipczuk et al., and to “open all boxes” for ensuring that each step of the construction can be implemented efficiently in the distributed setting.
Second, previous results consider model checking of first-order logics over arbitrary relational structures, and each relational structure has an associated graph (the Gaifman graph). Model checking over graphs of bounded expansion actually refers to model checking of relational structures whose Gaifman graphs are of bounded expansion. The techniques of Pilipczuk et al. (as well as the other related approaches) consist of enriching the structure relative to an arbitrary first-order formula with new symbols and new predicates for obtaining a quantifier-free formula. Enriching the structure naturally modifies the Gaifman graphs. In contrast, in distributed computing models such as CONGEST, nodes establish communications over a network that does not change. When enriching the structure, one thus need to be very careful for avoiding creating connections between nodes that cannot directly communicate in the network.
Given an FO formula $\varphi$, and a class $\mathcal{G}$ of labeled graphs with expansion $f$, quantifier elimination constructs a quantifier-free formula $\widehat{\varphi}$. This formula satisfies that, for each labeled graph $(G,L)\in\mathcal{G}$, there is a labeling $\widehat{L}$ such that $(G,L)\models\varphi\iff(G,\widehat{L})\models\widehat{\varphi}$. To build $\widehat{\varphi}$, specific types of unary and binary predicates are introduced. The new unary predicates, called label predicates, check the presence of certain symbols in the labels assigned to each node. The new binary predicates, called least common ancestor (or lca for short) predicates check the depth of the least common ancestor of two nodes in certain subtrees of the input graph.
@@ 1.2.1 Distributed Model Checking
The technique for building the pair $(\widehat{\varphi},\widehat{L})$ from $(\varphi,G,L)$ consists in first rewriting $\varphi$ so that it uses solely negations, disjunctions, and existential quantifiers. Under this format, the main issue becomes the elimination of the existential quantifiers. The technique of [71] proceeds gradually, by eliminating each existential quantifier one after the other. The basic step of the elimination is therefore the following. Let us assume that $\varphi(\bar{x})=\exists y\,\psi(\bar{x},y)$ with $\bar{x}=(x_{1},\dots,x_{\ell})$ is a vector of free variables (these variables are actually quantified in the overall formula, but the corresponding quantifiers will be eliminated at further steps). Let us moreover assume that the pair $(\widehat{\psi},\widehat{L}_{\psi})$ has already been computed from $(\psi,G,L)$, where $\widehat{\psi}$ is quantifier-free. The goal is then to build $(\widehat{\varphi},\widehat{L}_{\varphi})$ from $(\varphi,G,L)$ where $\widehat{\varphi}$ is quantifier-free, and $\widehat{L}_{\varphi}$ is a labeling obtained from $\widehat{L}_{\psi}$ by adding unary predicates.
@@
The existential quantifier in $\varphi(\bar{x})=\exists y\,\widehat{\psi}(\bar{x},y)$ is eliminated in two stages. First, recall that the characterization theorem from [66] states that, for any graph class $\mathcal{G}$ of expansion $f$, there exists a function $g:\mathbb{N}\to\mathbb{N}$ satisfying that, for every $k\geq 1$, the vertices of any graph $G\in\mathcal{G}$ can be colored using a palette of $g(k)$ colors such that, for any set $S$ of at most $k$ colors, the subgraph of $G$ induced by the vertices with colors in $S$ has treedepth at most $k$. By setting $k=|\varphi|$, i.e., by letting $k$ be the number of variables occurring in $\varphi$, it is guaranteed that at least one set $S$ of $k$ colors captures $k$ nodes for which $\varphi$ holds whenever such nodes exist. For instance, for $C_{4}$ freeness, for any coloring of the nodes with at least 4 colors, there exists a set $S$ of at most 4 colors such that, whenever there is a 4-cycle in the graph, this 4-cycle appears in the subgraph induced by the vertices colored with some color in $S$. Assuming given an appropriate $g(k)$-coloring, the quantifier elimination can proceed by considering all sets of $k$ colors $S$, and eliminating the existential quantifier in graphs of bounded treedepth. The second stage consists of eventually eliminating the existential quantifier by transforming a formula on labeled graphs of bounded treedepth, with an existential quantifier, into a formula on labeled forests of bounded depth without existential quantifier. This is achieved by encoding the existential quantifier into a disjunction of quantifier-free formulas, each referring to a specific finite subgraph pattern. The existence of the subgraph pattern is encoded in a complicated Boolean combination of label and lca predicates.
Our main contribution is to show that the whole process of quantifier elimination sketched above can be implemented in $\mathcal{O}(D+\log n)$ rounds in CONGEST, where the constant in the big-O notation depends solely on the formula $\varphi$ and the expansion $f$ of the graph class $\mathcal{G}$. Given $k=|\varphi|$, we use the distributed algorithm of [68] to compute the appropriate $g(k)$-coloring of the input graph $G$, in $\mathcal{O}(\log n)$ rounds. Granted with this coloring, we show how to transform $\varphi$ into $\widehat{\varphi}$, and to label $G$ for producing the convenient labeled graph $\widehat{L}$. For this purpose, we show how each node can compute the value of the unary predicates involved in computing $\widehat{L}$.
Gaifman’s Locality Theorem [35] states that every FO formula is equivalent to a Boolean combination of sentences saying that there exist vertices that are far from each other, each satisfying some local condition described by a FO formula referring to a fixed-size neighborhood of the vertex. Each local condition can be checked in each of the components of the considered subgraph of $G$ with bounded treedepth resulting from picking one set $S$ of $k$ colors, and the truth value of each lca and label predicate at every node can be computed in a constant number of rounds. The label predicates are also used for combining the results of each local condition. This combination involves nodes that are far away one from another in the graph, but we show that it can be done in $\mathcal{O}(D)$ rounds in CONGEST despite the limited bandwidth.
1.2.2 Counting and Optimization

### v2

@@ 1.2.1 Extension 1: Deciding FO formulas
Let $\Lambda$ be a finite set. For every FO formula $\varphi$ on $\Lambda$-labeled graphs, and for every class of graphs $\mathcal{G}$ of bounded expansion, there exists a distributed algorithm that, for every $n$-node network $G=(V,E)\in\mathcal{G}$ of diameter $D$, and every $\ell:V\to 2^{\Lambda}$, decides whether $(G,\ell)\models\varphi$ in $\mathcal{O}(D+\log n)$ rounds under the CONGEST model.
Note that bounded degeneracy would not suffice to establish the same result as above. Indeed, we show that the mere FO formula $C_{6}$-freeness requires $\tilde{\Omega}(\sqrt{n})$ rounds to be checked in CONGEST, even in graphs of degeneracy 2 and of diameter 6 (see Theorem 7.4). Moreover, our result is the best that one can hope regarding the class of formulas, in the sense that it cannot be extended to, e.g., MSO. Indeed, we show that non-3-colorability requires $\tilde{\Omega}(n)$ rounds to be checked in bounded degree graphs with logarithmic diameter (see Theorem 7.2).
1.2.2 Extension 2: Counting, Optimization, and Certification
Our results extend beyond model checking, in several directions. First, we consider the problem of counting solutions of FO formulas with free variables. We actually consider two forms of counting:
  • Local counting: Given a local formula $\varphi(x_{1},\dots,x_{k})$ with $k\geq 1$ free variables, such as $\textsf{adj}(x_{1},x_{2})\land\textsf{adj}(x_{2},x_{3})\land\textsf{adj}(x_{3},x_{1})$ (i.e., $x_{1},x_{2},x_{3}$ form a triangle), each node $v$ of the considered graph $G=(V,E)$ outputs a (possibly negative) integer $\nu(v)$ satisfying that the number of tuples $v_{1},\dots,v_{k}$ for which $(G,\ell)\models\varphi(v_{1},\dots,v_{k})$ equals $\sum_{v\in V}\nu(v)$.
  • Global counting: Given a general formula $\varphi(x_{1},\dots,x_{k})$ with $k\geq 1$ free variables, all nodes output a same value equal to the number of tuples $v_{1},\dots,v_{k}$ for which $(G,\ell)\models\varphi(v_{1},\dots,v_{k})$. E.g., counting the $k$-tuples of nodes of degree 3.
Theorem 1.3.
@@ 1.2.2 Extension 2: Counting, Optimization, and Certification
Let $\Lambda$ be a finite set. For every FO formula $\varphi(x_{1},\dots,x_{k})$ on $\Lambda$-labeled graphs with $k\geq 1$ free variables $x_{1},\dots,x_{k}$, and for every class of graphs $\mathcal{G}$ of bounded expansion, there exists a distributed algorithm that, for every $n$-node network $G=(V,E)\in\mathcal{G}$ of diameter $D$ and every $\ell:V\to 2^{\Lambda}$, performs global counting in ${O(D+\log n)}$ rounds in CONGEST. If the formula $\varphi(x_{1},\dots,x_{k})$ is local then there exists a distributed algorithm that performs local counting in $\mathcal{O}(\log n)$ rounds in CONGEST.
@@
Note that, in contrast to our Theorem 1.3, the best known algorithm for (locally) counting triangles in general graphs takes $\tilde{\mathcal{O}}(n^{1/3})$ rounds [9].
Some optimization problems can also be expressed as finding optimal solutions for FO formulas with free variables, over weighted graphs, e.g., finding a maximum-weight triangle. (If no triangles exist in the graph, then all nodes reject.) Note that, as usual in the literature on CONGEST [57], the weights assigned to the nodes or edges are assumed to be polynomial in the network size so that any weight can be transmitted between neighbors in a single round of communication. We prove the following theorem.
Theorem 1.4.
Let $\Lambda$ be a finite set. For every FO formula $\varphi(x_{1},\dots,x_{k})$ on $\Lambda$-labeled graphs with $k\geq 1$ free variables $x_{1},\dots,x_{k}$, and for every class graphs $\mathcal{G}$ of bounded expansion, there exists a distributed algorithm that, for every $n$-node network $G=(V,E)\in\mathcal{G}$ of diameter $D$, every $\ell:V\to 2^{\Lambda}$, and every weight function $\omega:V\to\mathbb{N}$, computes a $k$-tuple of vertices $(v_{1},\dots,v_{k})$ such that $(G,\ell)\models\varphi(v_{1},\dots,v_{k})$ and $\sum_{i=1}^{k}\omega(v_{i})$ is maximum, in $\mathcal{O}(D+\log n)$ rounds under the CONGEST model. (The same holds if replacing maximum by minimum.) If no tuples $(v_{1},\dots,v_{k})$ exist such that $(G,\ell)\models\varphi(v_{1},\dots,v_{k})$, then all nodes reject.
Theorem 1.4 allows to find, for example, a monochromatic triangle of minimum weight, or a multi-colored $k$-independent set of maximum weight, in a class $\mathcal{G}$ of $c$-colored graphs of bounded expansion.
Last but not least, our results extend to distributed certification. Roughly, a distributed certification scheme [36, 48] for a graph property $\varphi$ is a pair prover-verifier. The prover is a centralized, computationally unlimited oracle that assigns certificates to the nodes. The verifier is a 1-round distributed algorithm whose role is to check whether the certificates assigned to the nodes form a proof that the graph satisfies the property. A distributed certification scheme is correct if it satisfies completeness (i.e., if $(G,\ell)\models\varphi$, then the prover can assign certificates such that the verifier accepts at all nodes), and soundness (i.e., if $(G,\ell)\not\models\varphi$, then, no matter the certificates assigned to the nodes, the verifier rejects in at least one node). The main complexity measure is the size of the certificates assigned by the prover on legal instances.
Theorem 1.5.
@@
Let $\Lambda$ be a finite set. For every FO formula $\varphi$ on $\Lambda$-labeled graphs, and for every class of graphs $\mathcal{G}$ of bounded expansion, there exists a distributed certification scheme that, for every $n$-node network $G=(V,E)\in\mathcal{G}$ and every $\ell:V\rightarrow 2^{\Lambda}$, certifies $(G,\ell)\models\varphi$ using certificates on $\mathcal{O}(\log n)$ bits.
@@
Note that, with $O(\log n)$-bit certificates, the verifier can be implemented in a single round of communication in the CONGEST model, as desired. Note also that our certification scheme is actually a proof-labeling scheme [48], which is a more demanding form of distributed certification than, e.g., locally checkable proofs [36]. As a direct consequence of Theorem 1.5, certifying $C_{4}$-freeness can be done with $\mathcal{O}(\log n)$-bit certificates in graphs of bounded expansion. In contrast, certifying $C_{4}$-freeness requires certificates on $\tilde{\Omega}(\sqrt{n})$ bits in general graphs, using the reduction from [17]. More generally, [22] shows that every MSO formula can be certified with $O(\log n)$-bit certificates in graphs of bounded treedepth. Theorem 1.5 extends this result. Indeed, while Theorem 1.5 is for FO, and the result of [22] is for MSO, both logics have identical expressive power on graphs of bounded treedepth. Last but not least, as for our model-checking algorithm, Theorem 1.5 is the best that can be achieved. Indeed, we show that it does not extend to MSO, as certifying non-3-colorability requires certificates of $\tilde{\Omega}(\sqrt{n})$ bits in planar graphs (see Theorem 7.3), and even certificates of $\tilde{\Theta}(n)$ bits in bounded degree graphs (see Theorem 7.2).
Remark.
The open problem by Nešetřil and Ossona de Mendez was actually stated for the BROADCAST CONGEST model, i.e., a weaker version of the CONGEST model in which, at every round, each node $v$ must send the same $\mathcal{O}(\log n)$-bit message to all its neighbors (see, e.g.,  [17, 47]). However, (1) the problem is not necessarily easier to solve in the general version of the CONGEST model, and (2) we solve it anyway by the affirmative using an algorithm designed for the BROADCAST CONGEST model, and the complexity of our algorithm is expressed under the latter. For the sake of simplifying the notations, we simply refer to CONGEST in the statements of our theorems, but all our upper bound results (but Theorem 1.4) hold under the BROADCAST CONGEST model too [^3] The main reason why Theorem 1.4 does not hold in BROADCAST CONGEST is that our algorithm involves an information dissemination phase going downward a tree, in which different messages are sent to different children. .
@@ 1.3 Our Techniques
In a nutshell, our technical contribution is twofold:
  • The first contribution is an adaptation of the quantifier elimination method [18, 38, 59] from centralized computing to the distributed setting. On a general level, we primarily follow the steps of the centralized approach. Yet, this adaptation is highly non-trivial, as we had to perform all operations locally without incurring a high volume of communication between adjacent nodes.
  • The second, and perhaps main, contribution of the paper is the novel locality preserving quantifier elimination method. Specifically, given a local FO formula $\varphi(x)$, we show how to eliminate its quantifiers, i.e., producing a quantifier-free formula $\widehat{\varphi}(x)$ that is equivalent to $\varphi(x)$, without losing locality. This is achieved by generalizing the notion of locality to formulas with more than a single free variable in a specific way, thereby enabling the application of induction on the number of quantifiers while preserving locality.
More specifically, our approach to establish Theorem 1.1 is mainly based on the quantifier elimination methods from [59]. Recall that, given an FO formula $\varphi$, and a class $\mathcal{G}$ of graphs with expansion $f$, quantifier elimination transforms every FO formula $\varphi$ into a quantifier-free FO formula $\widehat{\varphi}$ such that, for every $G=(V,E)\in\mathcal{G}$, and every labeling $\ell:V\to 2^{\Lambda}$, one can construct a node-labeling $\widehat{\ell}$ of $G$ such that
$(G,\ell)\models\varphi\iff(G,\widehat{\ell})\models\widehat{\varphi}.$
In full generality, quantifier elimination allows the graph $G$ to be transformed into a graph $\widehat{G}$, with $(G,\ell)\models\varphi\iff(\widehat{G},\widehat{\ell})\models\widehat{\varphi}$. However, as we shall discuss later, the approach in [59] preserves the graph $G$. Yet, we have to overcome two main difficulties for computing $\widehat{\ell}$.
First, we have to implement quantifier elimination in the distributed setting. That is, we show how to compute the new labeling $\widehat{\ell}$ of $G$, and the new formula $\widehat{\varphi}$ efficiently in the CONGEST model. Note that for an algorithm to be efficient in CONGEST, not only should the exchange of information be restricted to nearby nodes in the graph, but also the amount of information exchanged between neighbors should be small. For instance, in $\mathcal{O}(\log n)$ rounds, a node with degree $\gg\log n$ cannot directly send the list of all its neighbors’ identifiers to any of its neighbors.
Second, we have to adapt quantifier elimination for preserving locality. Given $G,\ell$, and a local FO formula $\varphi$, we show how to compute $\widehat{\ell}$ and $\widehat{\varphi}$ so that not only $\widehat{\varphi}$ is quantifier-free, and satisfies, for every $v\in V$,
$(G,\ell)\models\varphi(v)\iff(G,\widehat{\ell})\models\widehat{\varphi}(v),$
but also every stage of the quantifier elimination process preserves locality. This is achieved, in particular, by extending the natural “semantic” notion of local formula $\varphi(x)$ with a unique free variable, to a “syntactic” notion of local formula $\psi(x,\overline{y})$ involving more than one free variables, i.e., $\bar{y}=(y_{1},\dots,y_{k})$ is a vector of free variables. This syntactical notion is based on relativization, which is, in our framework, defining a new structure by restricting its domain to elements satisfying a form of mutual proximity.
Let us now enter a bit more into the details.
1.3.1 Eliminating Existential Quantifiers
@@
In order to build the pair $(\widehat{\varphi},\widehat{\ell})$ from $(\varphi,G,\ell)$, the main issue is the elimination of the existential quantifiers (universal quantifiers can be handled as negated existential quantifiers). The technique of [59] proceeds gradually, by eliminating each existential quantifier one after the other. The basic step is therefore the elimination of one existential quantifier. Let us consider a formula $\psi(x,\bar{y})=\exists z\,\psi^{\prime}(x,\bar{y},z)$ where $\bar{y}=(y_{1},\dots,y_{k})$ is a vector of free variables — these variables are actually quantified in the original formula $\varphi(x)$, but the corresponding quantifiers will be eliminated at further steps. Let us assume that a pair $(\widehat{\psi^{\prime}},\ell^{\prime})$ has already been computed from $(\psi^{\prime},G,\ell)$, i.e., $\widehat{\psi^{\prime}}$ is quantifier-free, and $\ell^{\prime}$ is the appropriate new node-labeling of $G$. The goal is then to compute $(\widehat{\psi},\widehat{\ell})$ from $(\widehat{\psi^{\prime}},G,\ell^{\prime})$ where $\widehat{\psi}$ is quantifier-free, and $\widehat{\ell}$ is the appropriate new node-labeling. The existential quantifier in $\psi(x,\bar{y})=\exists z\,\widehat{\psi^{\prime}}(x,\bar{y},z)$ is eliminated in two stages.
@@
First, recall that the aforementioned characterization theorem from [53] states that, for any graph class $\mathcal{G}$ of expansion $f$, there exists a function $g:\mathbb{N}\to\mathbb{N}$ satisfying that, for every $k\geq 1$, the vertices of any graph $G\in\mathcal{G}$ can be colored using a palette of $g(k)$ colors such that, for any set $U$ of at most $k$ colors, the subgraph of $G$ induced by the vertices with colors in $U$ has treedepth at most $k$. By setting $k=|\psi|$, i.e., by letting $k$ be the number of variables occurring in $\psi$, it is guaranteed that at least one set $U$ of $k$ colors captures $k$ nodes for which $\psi$ holds whenever such nodes exist. For instance, for $C_{4}$ freeness, for any coloring of the nodes with at least 4 colors, there exists a set $U$ of at most 4 colors such that, whenever there is a 4-cycle in the graph, this 4-cycle appears in the subgraph induced by the vertices colored with some color in $U$. Assuming given an appropriate $g(k)$-coloring, the quantifier elimination can proceed by considering all sets $U$ of $k$ colors, and eliminating the existential quantifier in graphs of bounded treedepth.
The second stage involves eventually eliminating the existential quantifier by transforming a formula on labeled graphs of bounded treedepth with an existential quantifier into a formula on labeled forests of bounded depth without existential quantifiers. This is achieved by encoding the existential quantifier into a disjunction of quantifier-free formulas, each referring to a specific finite subgraph pattern. To encode the subgraph pattern, new unary and binary predicates are introduced. These unary predicates are called label predicates. Their aim is to specify the presence of certain symbols in the labels assigned to the nodes. The new binary predicates are called least common ancestor predicates, or simply lca predicate for short. Their aim is to specify the depth of the least common ancestor of two nodes in certain trees of the forest. The subgraph patterns are encoded as Boolean combinations of label and lca predicates.
1.3.2 Locality-Preserving Quantifier Elimination
A first novelty of this paper is to adapt the approach of [59] such that each step preserves the locality property of the considered formula. To that end, we extend the notion of local formula to more that one free variables: Roughly, a formula $\psi(x,\bar{y})$ is local if (1) $(G,\ell)\models\varphi(x,\bar{y})$ implies that $\bar{y}\subseteq B_{G}(x,r)$, i.e., each $y_{i}$ is at distance at most $r$ from $x$, and (2) for every two graphs $G$ and $G^{\prime}$, and every two vertices $x$ and $x^{\prime}$ of $G$ and $G^{\prime}$, respectively, if there is a (label preserving) isomorphism $h:B_{G}(x,r)\to B_{G^{\prime}}(x^{\prime},r)$ satisfying $h(x)=x^{\prime}$, then $(G,\ell)\models\varphi(x,\bar{y})\iff(G^{\prime},\ell)\models\varphi(h(x),h(\bar{y}))$. This “semantic” definition is difficult to use in practice, which is why we introduce a “syntactic” notion of local form for local formulas. For this purpose, we introduce new predicates, called distance predicates. These predicates are of the form $\mathrm{ball}(y,x,r)$, interpreted as “$y$ is in the ball of radius $r$ around $x$ in the graph”. The distance predicates allow us to syntactically enforce the locality property of a formula. For instance, in a formula $\mathrm{ball}(y,x,r)\wedge\varphi(x,y)$, the distance predicate enforces the fact that $x$ and $y$ are at bounded distance. At each step of the quantifier elimination process, we make sure that both the original formula $\varphi$, and the transformed formula $\widehat{\varphi}$ explicitly contain distance predicates that guarantee their locality.
The presence of negations in the formula, potentially induced by universal quantifiers treated as negated existential quantifiers, must be treated with care. Indeed, the locality property is not stable by negation. For instance, the negation of $\mathrm{ball}(y,x,r)\wedge\varphi(x,y)$ is $\neg\mathrm{ball}(y,x,r)\vee\neg\varphi(x,y)$, which is not local because $\neg\mathrm{ball}(y,x,r)$ is satisfied for nodes $y$ that are far away from $x$. To overcome this issue, we introduce redundancy in the syntax of local formulas. That is, instead of adding distance predicates once in the formula (for instance, immediately after the innermost quantifier), we duplicate the distance predicates wherever they are needed.
For instance, let us consider a formula $\varphi(x,\bar{y})=\forall z\,\psi(x,\bar{y},z)$ that is assumed to be local. As said before, quantifier elimination proceeds by eliminating existential quantifiers, and thus this formula is rewritten as $\varphi(x,\bar{y})=\neg\exists z\,\neg\psi(x,\bar{y},z)$. Now, we make the locality of $z$ explicit by rewriting $\varphi$ as
$\varphi(x,\bar{y})=\neg\exists z\,\big{(}\mathrm{ball}(z,x,r)\land\neg\psi(x,\bar{y},z)\big{)}.$
Moreover, we use redundancy, and reinforce the locality of $\bar{y}$ by introducing another distance predicate:
$\varphi(x,\bar{y})=\mathrm{ball}(\bar{y},x,r)\land\neg\exists z\,\big{(}\mathrm{ball}(z,x,r)\land\neg\psi(x,\bar{y},z)\big{)}.$
By distributing the distance predicate on $\bar{y}$, we get
$\varphi(x,\bar{y})=\mathrm{ball}(\bar{y},x,r)\land\neg\exists z\,\big{(}\mathrm{ball}(\bar{y},x,r)\land\mathrm{ball}(z,x,r)\land\neg\psi(x,\bar{y},z)\big{)},$
which can be rewritten as
$\varphi(x,\bar{y})=\mathrm{ball}(\bar{y},x,r)\land\neg\exists z\,\big{(}\mathrm{ball}((\bar{y},z),x,r)\land\neg\psi(x,\bar{y},z)\big{)}.$
As a result, the potential loss of locality resulting from the negation of $\psi(x,\bar{y},z)$ is recovered thanks to the (redundant) distance predicate, as $\mathrm{ball}((\bar{y},z),x,r)\land\neg\psi(x,\bar{y},z)$ is local. Eliminating the existential quantifier in this formula yields a (local) quantifier-free formula. Negation is, however, applied a second time, which may result in losing locality. Yet, locality is again recovered using the other (redundant) distance predicate $\mathrm{ball}(\bar{y},x,r)$.
1.3.3 Quantifier Elimination in CONGEST
The second novelty of this paper is the implementation of quantifier elimination, including the new constraint of preserving locality, in CONGEST, specifically in BROADCAST CONGEST. We had to tackle two main difficulties.
First, the distributed implementation of the techniques in [59] requires technical attention in several places. There are indeed operations that are very simple to perform in the centralized setting (they are even often implicit when described in the centralized algorithm), but it is not necessarily clear whether all these operations are local (i.e., involving nodes that are close in the graph), and frugal in term of bandwidth (i.e., requiring neighboring nodes to exchange a limited amount of information). As a consequence, we had to carefully revisit the entire construction and ’open all boxes’ to ensure that each step of the construction can be implemented efficiently in (BROADCAST) CONGEST.
Second, model checking over graphs of bounded expansion actually refers to model checking of relational structures whose Gaifman graphs are of bounded expansion. The techniques from [59] (as well as other related approaches) involve enriching the structure relative to an arbitrary first-order formula with new symbols and predicates to obtain a quantifier-free formula. Enriching the structure naturally modifies Gaifman graphs. In contrast, in distributed computing models such as CONGEST, nodes establish communications over a network that does not change. When enriching the structure, one must therefore be very careful to avoid creating connections between nodes that cannot directly communicate in the network.
1.3.4 Model Checking, Counting, Optimization, and Certification
Model Checking.
For checking whether $(G,\ell)\models\varphi$ in CONGEST, for a given FO formula $\varphi$, and a given labeled graph $G\in\mathcal{G}$, we use the same techniques as the ones used to mark all nodes satisfying a given local FO formula. That is, the proof of Theorem 1.2 follows the same guidelines as the ones used to establish Theorem 1.1. On the one hand, the approach is simplified as one does not need to care about relativization. On the other hand, the ”non-locality” of the formula requires checking properties that potentially occur at nodes that are far away from each other in the graph (e.g., the presence of at least two nodes of degree 3). In the context of quantifier elimination, this translates into counting the number of nodes in which unary predicates are satisfied. This counting is merely achieved along the edges of a breadth-first search (BFS) tree, in $\mathcal{O}(D)$ rounds. We show that all other operations involved in quantifier elimination can be performed in $\mathcal{O}(\log n)$ rounds, for a total of $\mathcal{O}(D+\log n)$ rounds.
Counting and Optimization.
Establishing Theorem 1.3, i.e., showing how to efficiently count the number of $k$-tuples of vertices $(v_{1},\dots,v_{k})$ such that $(G,\ell)\models\varphi(v_{1},\dots,v_{k})$, requires more work. For the global counting problem, we proceed in two stages.